### v1

@@ 5 Experimental Results
[…]
@@ 5.2 PoseSLAM optimization results
The proposed algorithm was tested on the four different scenarios using the following parameters: a translation threshold of $\delta t_{\text{trans}}=1.0$ meters, a rotation threshold of $\delta q_{\text{rot}}=5^{\circ}$, and constraints were excluded if they had a fitness score $f>3.5$. The weight matrices were set as follows: $w_{o}$ was kept as $[1.0,1.0,1.0,1.0]$, without influencing the weight of the residual, while $w_{p}$ was defined as $[0.1,0.1,0.1]$ to scale the components based on the vehicle’s orientation.
The results of the optimization are presented in Fig. 4, where both top and side views are shown for a comprehensive comparison of trajectories. Table 2 provides metrics for both direct and optimized odometry along with the variations of 4DRadarSLAM [13] (GICP and ADPGICP -which incorporates a probabilistic distribution into the GICP algorithm-, with and without loop closure). In our case, our approach does not include large loop closure, but it may be of interest to compare all metrics to assess how far they are from the loop closure algorithm also proposed in [13].
When comparing our optimized trajectory with the direct one, it is evident from the results that the optimization notably reduces the drift in alignment with the ground truth for most maps. An exception is the “loop1” map, where the algorithm proves less effective at mitigating drift, likely due to noisier sensor readings in this environment coupled with a trajectory containing more curves.
Comparison with previous results from [26] shows similar performance levels. We focus primarily on absolute error, as it is the more critical metric, though relative errors of similar magnitude are also provided. The analyses indicate clear improvements in absolute error across most maps, except in “loop1,” where the error reduction is minimal, though it remains within a comparable range. We can also see that the relative errors (translational and rotational) are very similar to those [26], while we improve, in general, in absolute errors. It should be noted that loop closure algorithms are shown here for reference but are not directly compared, as they entail an additional optimization step not implemented in pure odometric methods. That said, our approach also provides absolute errors very similar to the loop closure methods.
@@
Velocity (ours) 5.570 0.129 1.934 18.069 0.1040 103.518 14.471 0.018 288.696 15.233 0.023 401.027
SLAM (ours) 2.033 0.048 1.181 3.989 0.0190 11.582 6.779 0.007 144.837 5.128 0.007 45.84
4DRadarSLAM gicp* 4.13 0.055 3.96 4.62 0.0184 17.42 4.84 0.0060 132.92 3.22 0.0060 57.29
4DRadarSLAM adpgicp* 3.56 0.0369 2.61 3.55 0.0171 21.30 6.09 0.0082 227.54 4.09 0.0097 59.12
4DRadarSLAM gicp-lc* 2.79 0.0511 2.54 3.10 0.0120 14.34 4.12 0.0065 68.88 3.46 0.0052 72.28
4DRadarSLAM adpgicp-lc* 3.02 0.0448 2.35 2.85 0.0131 11.37 5.79 0.0100 84.88 4.03 0.0069 43.67
Table 2: Comparison of relative translation ($t_{\text{rel}}$), relative rotation ($r_{\text{rel}}$), and absolute translation ($t_{\text{abs}}$) errors across different datasets and algorithms. The results for *-lc algorithms are shown for informational purposes only.
@@ 6 Conclusions
As future work, a tightly coupled methodology between the radar and IMU is intended, not only allowing the IMU to influence radar data but also enabling radar information to enhance IMU orientation estimation. The aim is to improve the IMU preintegration system by enabling the estimation of parameters such as gyroscope and accelerometer bias through additional information provided to the implemented filter, thus enhancing the accuracy of the system’s orientation evolution.
@@
Loop closure systems for radar data must address with challenges such as the noise and sparsity of radar measurements, particularly affecting the vertical component, as discussed in this report. Addressing these challenges is crucial for enhancing the robustness of radar-based odometry, allowing it to better handle long-term drift and improve the overall accuracy of the system.
@@ Acknowledgements
This work was supported by the grant INSERTION PID2021-127648OB-C31, funded by MCIN/AEI/ 10.13039/501100011033 and “European Union NextGenerationEU/PRTR”.

### v2

@@ 6 Experimental Results
To evaluate the performance of our algorithm, we used the rpg-trajectory-evaluation package [23], which is the same tool employed by the radar SLAM algorithm [13], with which we will later compare. This package provides key metrics such as RMSE (Root Mean Square Error) values for relative translation $t_{\text{rel}}(\%)$, relative rotation $r_{\text{rel}}(\text{deg/m})$, and absolute translation $t_{\text{abs}}(\text{m})$.
@@ 6.2 PoseSLAM optimization results
The proposed algorithm was tested on the four different scenarios using the following parameters: a translation threshold of $\delta t_{\text{trans}}=1.0$ meters, a rotation threshold of $\delta q_{\text{rot}}=5^{\circ}$, and a fitness score threshold of $f_{\text{th}}=3.5$. The weight matrices were set as follows: $w_{o}$ was kept as $[1.0,1.0,1.0,1.0]$, without influencing the weight of the residual, while $w_{p}$ was defined as $[0.1,0.1,0.1]$ to scale the components based on the vehicle’s orientation.
The results of the optimization are presented in Fig. 6, where both top and side views are shown for a comprehensive comparison of trajectories. Table 2 provides metrics for both direct and optimized odometry along with the variations of 4DRadarSLAM [13] (GICP and ADPGICP, which incorporates a probabilistic distribution into the GICP algorithm, with and without loop closure). We focus primarily on absolute error, highlighted in bold in Table 2, as it is the most critical metric, though relative errors are also provided. In our case, our approach does not include large loop closure, but it may be of interest to compare all metrics to assess how far they are from the loop closure algorithm also proposed in [13].
When comparing our optimized trajectory with the direct one, it is evident from the results that the optimization notably reduces the drift in alignment with the ground truth. Comparison with previous results from [13] shows that our method achieves improved performance, with clear reductions in absolute error across all maps, while relative translational and rotational errors remain similar. It should be noted that loop closure algorithms are shown here for reference but are not directly compared, as they entail an additional optimization step not implemented in pure odometric methods. That said, our approach also provides absolute errors very similar to the loop closure methods, and even outperforms them on certain trajectories.
@@
Method 1 5.570 0.129 1.934 18.069 0.1040 103.518 14.471 0.018 288.696 15.233 0.023 401.027
Method 2 1.871 0.038 0.988 2.489 0.0111 6.414 5.447 0.006 104.959 4.549 0.006 52.976
Method 3 4.13 0.055 3.96 4.62 0.0184 17.42 4.84 0.0060 132.92 3.22 0.0060 57.29
Method 4 3.56 0.0369 2.61 3.55 0.0171 21.30 6.09 0.0082 227.54 4.09 0.0097 59.12
Method 5 2.79 0.0511 2.54 3.10 0.0120 14.34 4.12 0.0065 68.88 3.46 0.0052 72.28
Method 6 3.02 0.0448 2.35 2.85 0.0131 11.37 5.79 0.0100 84.88 4.03 0.0069 43.67
Table 2: Comparison of relative translation ($t_{\text{rel}}$), relative rotation ($r_{\text{rel}}$), and absolute translation ($t_{\text{abs}}$) errors across different datasets and algorithms. Method 1: Velocity (ours), Method 2: SLAM (ours), Method 3: 4DRadarSLAM gicp, Method 4: 4DRadarSLAM adpgicp, Method 5: 4DRadarSLAM gicp-lc, Method 6: 4DRadarSLAM adpgicp-lc. The results for -lc algorithms are shown for informational purposes only.
6.3 Runtime and memory consumption
The experiments were performed on an HP Victus 16 laptop equipped with 32GB RAM and a 13th generation Intel Core i7-13700H processor. The software implementation was divided into two different processes: the Cloud Processing Node in charge of the point-cloud filtering and ego-velocity computation, and the Optimizer Node dealing with the graph computation, update and optimization. The complete system maintains a moderate memory footprint, with CPU usage consistently reaching 4.2% for the Optimizer Node and around 0.9% for the Cloud Processing Node. Memory consumption remains stable at approximately 49.5 MB (0.15 % of total system memory) for the optimizer and 11.2MB (0.03% of total system memory) for the cloud processing.
Regarding runtime, the system has been tested across several datasets, with execution times generally following similar patterns for both nodes. Thus, the Cloud Processing Node operates at an average execution time of $8.91\pm 2.27$ miliseconds (see Table 3). More specifically, Pre-Processing accounts for $68.3\%$ of the total execution time, Ego-Velocity Estimation takes up $28.36\%$, and Cloud Filtering represents $3.26\%$.
The Optimizer Node has a more complex execution pattern, as it depends on whether a new keyframe is added (see Table 4). When no keyframe is added, the process involves only Direct Odometry Estimation, which takes an average of $0.12\pm 0.06$ miliseconds, resulting in minimal computational overhead. However, when a new keyframe is added, the optimization process significantly increases the execution time to an average of $713.52\pm 334.10$ miliseconds. This is primarily due to the Update Constraints process, which accounts for $99.86\%$ of the total time. The remaining computational cost is distributed among Direct Odometry Estimation ($0.04\%$), Graph Creation ($0.02\%$), and Graph Optimization ($0.07\%$).
Although optimization introduces a delay exceeding the radar’s operational frequency of $10$ Hz, this impact is mitigated by the fact that keyframe additions occur less frequently than direct odometry updates. As a result, the overall runtime of the optimizer node averages $29.70\pm 157.00$ miliseconds, ensuring that the delay is balanced across multiple iterations.
[…]
Table 3: Execution time analysis of the Cloud Processing Node
[…]
Table 4: Execution time analysis of the Optimizer Node
@@ 7 Conclusions
As future work, a tightly coupled approach between the radar and IMU is intended, not only allowing the IMU to influence radar data but also enabling radar-derived egovelocity to enhance IMU orientation estimation. The aim is to improve the IMU preintegration system by allowing the estimation of parameters such as gyroscope and accelerometer bias through additional information provided to the implemented filter, thus enhancing the accuracy of the system’s orientation evolution.
@@
Loop closure systems for radar data must address challenges such as the noise and sparsity of radar measurements, particularly affecting the vertical component, as discussed in this report. Addressing these challenges is crucial for enhancing the robustness of radar-based odometry, allowing it to better handle long-term drift and improve the overall accuracy of the system.
@@ Acknowledgements
Funding for open access publishing: Universidad Pablo de Olavide/CBUA. This work was partially supported by the grants INSERTION (PID2021-127648OB-C31) and NORDIC (TED2021-132476B-I00), both funded by the “Agencia Estatal de Investigación – Ministerio de Ciencia, Innovación y Universidades” and the “European Union NextGenerationEU/PRTR”.